OpenStudy (anonymous):

'A' alone can complete a piece of work in 8 days .work done by B alone in one day is half of the work done by A alone in one day.in how many days the work gets completed when they work together

OpenStudy (anonymous):

If the difference between the simple interest and com�pound interest earned on an amount @ 15 p.c.p.a. at the end of 3 years is Rs 595.35, what is the amount

OpenStudy (anonymous):

let the principal be p SI = PRT /100 = px15x3/100 = 45p/100 = 9p/20 CI = A-P = P(1+ r/100)^n-p = p[(1+15/100)^3-1] now ur CI-SI = 595.35 solve and you will get the answer

OpenStudy (anonymous):

i got the same answer

OpenStudy (anonymous):

k @strgreen

OpenStudy (anonymous):

595.35